# Non-Convex Stochastic Composite Optimization with Polyak Momentum

### v1

@@ Appendix C Missing proofs in Section 7
$-R_{k}\leq-\frac{4{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{17(M^{2}+L^{2})}+\frac{16\Delta_{k}}{17(M^{2}+L^{2})}+\frac{16S_{k}}{17(M^{2}+L^{2})}$ (11)
Recall in the proof of Lemma 5.6, we have:
$\Phi_{k+1}$ $\leq F_{k}-\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)R_{k}+a\gamma^{2}\sigma^{2}$
$\quad+\left(1-\gamma+\frac{1}{a(M-L)}\right)a\Delta_{k}$
If $\frac{M-L}{4}-\frac{aL^{2}}{\gamma}>0$, then we can plug Equation 11 in and get:
$\Phi_{k+1}$ $\leq F_{k}-\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)\left(\frac{4{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{17(M^{2}+L^{2})}-\frac{16\Delta_{k}}{17(M^{2}+L^{2})}-\frac{16S_{k}}{17(M^{2}+L^{2})}\right)$
$\quad+\left(1-\gamma+\frac{1}{a(M-L)}\right)a\Delta_{k}+a\gamma^{2}\sigma^{2}$
$=F_{k}-\frac{4}{17}\left(\frac{M-L}{4(M^{2}+L^{2})}-\frac{aL^{2}}{\gamma(M^{2}+L^{2})}\right){\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]$
$\quad+\left(1-\gamma+\frac{1}{a(M-L)}+\frac{16}{17}\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)\frac{1}{a(M^{2}+L^{2})}\right)a\Delta_{k}+a\gamma^{2}\sigma^{2}$
$\quad+\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)\frac{16S_{k}}{17(M^{2}+L^{2})}$
Now let
$\frac{M-L}{8(M^{2}+L^{2})}=\frac{aL^{2}}{\gamma(M^{2}+L^{2})}$
then $a=\frac{\gamma(M-L)}{8L^{2}}$. Note that with this $a$, we have $\frac{M-L}{4}-\frac{aL^{2}}{\gamma}>0$. Now
$\frac{1}{a(M-L)}+\frac{16}{17}\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)\frac{1}{a(M^{2}+L^{2})}$ $=\frac{8L^{2}}{\gamma(M-L)^{2}}+\frac{2(M-L)}{17}\frac{1}{a(M^{2}+L^{2})}$
$=\frac{8L^{2}}{\gamma(M-L)^{2}}+\frac{16L^{2}}{17\gamma(M^{2}+L^{2})}$
$=\frac{8L^{2}}{\gamma}\left(\frac{1}{(M-L)^{2}}+\frac{2}{17(M^{2}+L^{2})}\right)$
Therefore we need $\frac{8L^{2}}{\gamma}\left(\frac{1}{(M-L)^{2}}+\frac{2}{17(M^{2}+L^{2})}\right)\leq\gamma$, rearranging, it suffices to have:
$\gamma^{2}\geq 8L^{2}\left(\frac{1}{(M-L)^{2}}+\frac{2}{17(M^{2}+L^{2})}\right)$
Note that $(M-L)^{2}\leq M^{2}+L^{2}$, it suffices to set $\gamma^{2}=\frac{152}{17}\frac{L^{2}}{(M-L)^{2}}$, or equivalently:
$\gamma=\sqrt{\frac{152}{17}}\frac{L}{M-L}$
Note that this requires $M>\frac{(\sqrt{152}+\sqrt{17})L}{\sqrt{17}}$, where it suffices that $M>4L$. Now
$a=\frac{M-L}{8L^{2}}\sqrt{\frac{152}{17}}\frac{L}{M-L}=\sqrt{\frac{19}{156}}\frac{1}{L}$
Putting them all together,
$\Phi_{k+1}\leq\Phi_{k}-\frac{M-L}{34(M^{2}+L^{2})}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{4L}{(M-L)^{2}}\sigma^{2}+\frac{2(M-L)S_{k}}{17(M^{2}+L^{2})}$
Let $M=\tau L$ for some $\tau>4$, then
$\Phi_{k+1}\leq\Phi_{k}-\frac{\tau-1}{34(\tau^{2}+1)L}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{4}{(\tau-1)^{2}L}\sigma^{2}+\frac{2(\tau-1)}{17(\tau^{2}+1)L}S_{k}$
Note that for $\tau>4,\frac{\tau-1}{\tau^{2}+1}\geq\frac{1}{2\tau},\frac{1}{(\tau-1)^{2}}\leq\frac{2}{\tau^{2}}$ and $\frac{\tau-1}{\tau^{2}+1}\leq\frac{1}{\tau}$. Therefore, for $M>7L$:
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{68M}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{8L}{M^{2}}\sigma^{2}+\frac{2}{17M}S_{k}$
@@
Rearranging and summing Equation 10 over $k=0,\ldots,K-1$, we get:
@@
$\frac{1}{68M}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\Phi_{0}+\frac{8KL}{M^{2}}\sigma^{2}+\frac{2}{17M}\sum_{k=0}^{K-1}S_{k}$
@@
$\frac{1}{K}\sum_{k=0}^{K=1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\frac{68M\Phi_{0}}{K}+\frac{544L\sigma^{2}}{M}+\frac{8}{K}\sum_{k=0}^{K-1}S_{k}$
Recall that $M>4L$, then for $M=4L+\sqrt{\frac{8KL\sigma^{2}}{\Phi_{0}}}$, we have:
$\frac{1}{K}\sum_{k=0}^{K=1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\mathcal{O}\left(\sqrt{\frac{L\Phi_{0}\sigma^{2}}{K}}+\frac{L\sigma^{2}}{K}\right)+\frac{8}{K}\sum_{k=0}^{K-1}S_{k}$
Appendix D Proximal gradient mapping
In this section, we briefly comment on the proximal gradient mapping, which is the convergence criteria used in most existing works [16, 14, 39, 17, 37, 42]. Proximal gradient mapping is defined in a very general context, where we consider the constrained composite optimization problem:
$\min_{\mathbf{x}\in X}F(\mathbf{x})\coloneqq f(\mathbf{x})+\psi(\mathbf{x}),$
@@ Appendix D Proximal gradient mapping
and in the literature, the convergence is studied in terms of $\left\lVert P_{X}(\mathbf{x},\nabla f(\mathbf{x}),M)\right\rVert^{2}$. We point out that it is very easy to prove an analogous version of Lemma 5.3 and Lemma 5.5 in terms of $\left\lVert P_{X}(\mathbf{x},\nabla f(\mathbf{x}),M)\right\rVert^{2}$ (which implies that our results extend to the proximal gradient mapping case), but we omit the details here. Instead, we argue that the size of $\nabla F(\mathbf{x})$ is a more natural convergence criterion. Consider the simplest situation where $\psi$ is convex and differentiable, $X=\mathbb{R}^{d}$, and the mirror map $r(\mathbf{x})=\frac{1}{2}\left\lVert\mathbf{x}\right\rVert^{2}$ (i.e. the Euclidean geometry). Then we have the usual stationarity condition for $\mathbf{x}^{+}$:

### v4

@@ Appendix D Missing Proofs in Section 7
$\frac{17}{16}(M_{k}^{2}+L^{2})R_{k}\geq\frac{1}{4}G_{k+1}^{2}-\Delta_{k}-S_{k}.$ (12)
Recall in the proof of Lemma 5, we have:
$\Phi_{k+1}\leq F_{k}-H_{k}R_{k}+\left(1-\gamma_{k}+\frac{1}{a(M_{k}-L)}\right)a\Delta_{k}+a\gamma_{k}^{2}\sigma^{2},$
where $H_{k}\coloneqq\frac{M_{k}-L}{4}-\frac{aL^{2}}{\gamma_{k}}$. If $H_{k}>0$, then we can plug Equation 12 in and get:
$\Phi_{k+1}$ $\leq F_{k}-\frac{16H_{k}}{17(M_{k}^{2}+L^{2})}\left(\frac{1}{4}G_{k+1}^{2}-\Delta_{k}-S_{k}\right)+\left(1-\gamma_{k}+\frac{1}{a(M_{k}-L)}\right)a\Delta_{k}+a\gamma_{k}^{2}\sigma^{2}$
$=F_{k}-\frac{4H_{k}}{17(M_{k}^{2}+L^{2})}G_{k+1}^{2}+\left(1-\gamma_{k}+\frac{1}{a(M_{k}-L)}+\frac{16H_{k}}{17a(M_{k}^{2}+L^{2})}\right)a\Delta_{k}+a\gamma_{k}^{2}\sigma^{2}+\frac{16H_{k}S_{k}}{17(M_{k}^{2}+L^{2})}.$
Now let $a=\frac{\gamma_{k}(M_{k}-L)}{8L^{2}}$, so that $H_{k}=\frac{M_{k}-L}{8}$. Then,
$\frac{1}{a(M_{k}-L)}+\frac{16H_{k}}{17a(M_{k}^{2}+L^{2})}=\frac{8L^{2}}{\gamma_{k}(M_{k}-L)^{2}}+\frac{16L^{2}}{17\gamma_{k}(M_{k}^{2}+L^{2})}=\frac{8L^{2}}{\gamma_{k}}\left(\frac{1}{(M_{k}-L)^{2}}+\frac{2}{17(M_{k}^{2}+L^{2})}\right).$
Therefore we need $\frac{8L^{2}}{\gamma_{k}}\bigl{(}\frac{1}{(M_{k}-L)^{2}}+\frac{2}{17(M_{k}^{2}+L^{2})}\bigr{)}\leq\gamma_{k}$ or, equivalently, $\gamma_{k}^{2}\geq 8L^{2}\bigl{(}\frac{1}{(M_{k}-L)^{2}}+\frac{2}{17(M_{k}^{2}+L^{2})}\bigr{)}$. Since $(M_{k}-L)^{2}\leq M_{k}^{2}+L^{2}$, it suffices to set $\gamma_{k}^{2}=\frac{152}{17}\frac{L^{2}}{(M_{k}-L)^{2}}$, i.e., $\gamma_{k}=\sqrt{\frac{152}{17}}\frac{L}{M_{k}-L}$. Note that this requires $M_{k}>\frac{(\sqrt{152}+\sqrt{17})L}{\sqrt{17}}$, which is ensured whenever $M_{k}>4L$. For our choice of $\gamma_{k}$, we get $a=\frac{1}{8L}\sqrt{\frac{152}{17}}=\sqrt{\frac{19}{156}}\frac{1}{L}$. Putting everything together, we get
$\Phi_{k+1}\leq\Phi_{k}-\frac{M_{k}-L}{34(M_{k}^{2}+L^{2})}G_{k+1}^{2}+\frac{4L}{(M_{k}-L)^{2}}\sigma^{2}+\frac{2(M_{k}-L)S_{k}}{17(M_{k}^{2}+L^{2})}.$
Let $M_{k}=\tau_{k}L$ for some $\tau_{k}>7$. Note that $\frac{\tau_{k}-1}{\tau_{k}^{2}+1}\geq\frac{1}{2\tau_{k}}$, $\frac{1}{(\tau_{k}-1)^{2}}\leq\frac{2}{\tau_{k}^{2}}$ and $\frac{\tau_{k}-1}{\tau_{k}^{2}+1}\leq\frac{1}{\tau_{k}}$. Therefore, for $M_{k}>7L$, we get
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{68M_{k}}G_{k+1}^{2}+\frac{8L}{M_{k}^{2}}\sigma^{2}+\frac{2S_{k}}{17M_{k}}.\qed$
@@
Setting $M_{k}=M$ some constant, and rearranging and summing Equation 11 over $k=0,\ldots,K-1$, we get:
@@
$\frac{1}{68M}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\Phi_{0}+\frac{8KL}{M^{2}}\sigma^{2}+\frac{2}{17M}\sum_{k=0}^{K-1}S_{k}.$
@@
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\frac{68M\Phi_{0}}{K}+\frac{544L\sigma^{2}}{M}+\frac{8}{K}\sum_{k=0}^{K-1}S_{k}.$
Then for $M=7L+\sqrt{\frac{8KL\sigma^{2}}{\Phi_{0}}}$, we have:
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\mathcal{O}\biggl{(}\sqrt{\frac{L\Phi_{0}\sigma^{2}}{K}}+\frac{L\sigma^{2}}{K}\biggr{)}+\frac{8}{K}\sum_{k=0}^{K-1}S_{k}.$
Appendix E Sampling from a Stream of Data
Following Theorem 1, we briefly mentioned that one can efficiently sample the desired output point. In this section, we explain how to perform such sampling at no extra computation and memory cost. This might have been discussed in the literature, but we still provide a detailed explanation for completeness and the reader’s convenience.
Proposition 3.
Given a stream of points $\{\mathbf{x}_{k}\}_{k=1}^{\infty}$ in $\mathbb{R}^{d}$ and positive scalars $\{h_{k}\}_{k=1}^{\infty}$, we can maintain, at each step $k\geq 1$, the random variable $\mathbf{x}_{t(k)}$, where $t(k)$ is a random index from $\{1,\ldots,k\}$ chosen with probabilities $\Pr(t(k)=i)=\frac{h_{i}}{H_{k}}$, $i=1,\ldots,k$, where $H_{k}\coloneqq\sum_{i=1}^{k}h_{i}$. This requires only $\mathcal{O}(d)$ memory and computation.
Proof.
We maintain the variables $\hat{\mathbf{x}}_{k}\in\mathbb{R}^{d}$ and $H_{k}\in\mathbb{R}$ which are both initialized to 00 at step $k=0$. Then, at each step $k\geq 1$, we update $H_{k}\leftarrow H_{k-1}+h_{k}$ and also, with probability $\frac{h_{k}}{H_{k}}$, we update $\hat{\mathbf{x}}_{k}\leftarrow\mathbf{x}_{k}$ (or, with probability $1-\frac{h_{k}}{H_{k}}$, keep the old $\hat{\mathbf{x}}_{k}=\hat{\mathbf{x}}_{k-1}$). The memory and computation costs are $\mathcal{O}(d)$. Note that, for any $1\leq i\leq k$, the event $\hat{\mathbf{x}}_{k}=\mathbf{x}_{i}$ happens iff $\hat{\mathbf{x}}$ was updated at step $i$ and then not updated at each step $j=i+1,\ldots,k$. Hence, for any $1\leq i\leq k$, we have
$\Pr(\hat{\mathbf{x}}_{k}=\mathbf{x}_{i})=\frac{h_{i}}{H_{i}}\cdot\prod_{j=i+1}^{k}\left(1-\frac{h_{j}}{H_{j}}\right)=\frac{h_{i}}{H_{i}}\cdot\prod_{j=i+1}^{k}\frac{H_{j-1}}{H_{j}}=\frac{h_{i}}{H_{k}}.\qed$
Appendix F Convergence in the Non-Differentiable Case
In this section we briefly discuss the convergence of Algorithm 1 in the non-differentiable case. In this case we assume that $\psi$ is convex. We assume that each subproblem is solved exactly, i.e. $\mathbf{x}_{k+1}=\operatorname*{arg\,min}_{\mathbf{x}}\Omega_{k}(\mathbf{x})$. In particular, this implies that there exists subgradient $\mathbf{g}_{k+1}^{\psi}\in\partial\psi(\mathbf{x}_{k+1})$ such that:
$\mathbf{m}_{k}+\mathbf{g}_{k+1}^{\psi}+M(\mathbf{x}_{k+1}-\mathbf{x}_{k})=\mathbf{0}$
Therefore, for convenience, we define $\nabla\psi(\mathbf{x}_{k+1})\coloneqq\mathbf{g}_{k+1}^{\psi}$ as the specific subgradient of $\psi$ at $\mathbf{x}_{k+1}$ that we use. Similarly, we define $\nabla F(\mathbf{x}_{k+1})\coloneqq\nabla f(\mathbf{x}_{k+1})+\mathbf{g}_{k+1}^{\psi}=\nabla f(\mathbf{x}_{k+1})+\nabla\psi(\mathbf{x}_{k+1})$ as the specific subgradient of $F$ at $\mathbf{x}_{k+1}$ that we use. This way, all our previous proofs still hold, and we have the following:
Theorem 4.
Under Assumptions 1, 3 and 4, for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, if $M_{k}=M$ some constant, $\gamma\coloneqq\frac{3L}{M-L}$, there exists $M=4L+\frac{3^{\nicefrac{{3}}{{2}}}}{2}\sqrt{\frac{KL\sigma^{2}}{\Phi_{0}}}$ such that $K=\mathcal{O}\bigl{(}\frac{L\Phi_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{L\Phi_{0}}{\varepsilon}\bigr{)}$ iterations of Algorithm 1 is sufficient to achieve ${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{t})\right\rVert^{2}\right]\leq\varepsilon$, where $t$ is chosen from $\{1,\ldots,K\}$ uniformly at random.
We remark that this can be reformulated in terms of the distance between $\partial F$ and $\mathbf{0}$: we have ${\rm dist}^{2}(\partial F(\mathbf{x}_{t}),0)\leq\varepsilon$.
Appendix G Convergence Criterias
In this section, we discuss the differences and connections between the convergence criterias used in the literature.
Proximal Gradient Mapping:
The most popular convergenec criteria used in most of the earlier works on non-convex composite optimization is proximal gradient mapping Ghadimi et al. [2016], Ghadimi and Lan [2013a], Wang et al. [2019], Hendrikx et al. [2020], Tran-Dinh et al. [2022], Xu and Xu [2022]. Proximal gradient mapping is defined in a very general context, where we consider the constrained composite optimization problem:
$\min_{\mathbf{x}\in X}\bigl{[}F(\mathbf{x})\coloneqq f(\mathbf{x})+\psi(\mathbf{x})\bigr{]},$
@@ Appendix G Convergence Criterias
and in the literature, the convergence is studied in terms of $\left\lVert P_{X}(\mathbf{x},\nabla f(\mathbf{x}),M)\right\rVert^{2}$. We point out that it is very easy to prove analogous versions of lemmas 2 and 4 in terms of $\left\lVert P_{X}(\mathbf{x},\nabla f(\mathbf{x}),M)\right\rVert^{2}$ (which implies that our results extend to the proximal gradient mapping case), but we omit the details here. Instead, we argue that the size of $\nabla F(\mathbf{x})$ is a more natural convergence criterion. Consider the simplest situation where $\psi$ is convex and differentiable, $X=\mathbb{R}^{d}$, and the mirror map $r(\mathbf{x})=\frac{1}{2}\left\lVert\mathbf{x}\right\rVert^{2}$ (i.e. the Euclidean geometry). Then we have the usual stationarity condition for $\mathbf{x}^{+}$:
@@
Moreau Envelope:
A closely related convergence criterion that was proposed to address the non-convergence problem of proximal gradient mapping of the earlier works is the Moreau envelope [Davis and Drusvyatskiy, 2019]. In Euclidean geometry, for some parameter $\lambda$, the Moreau envelope is defined as the following:
$F_{\lambda}(\mathbf{x})\coloneqq\min_{\mathbf{y}}\Bigl{\{}F(\mathbf{y})+\frac{1}{2\lambda}\left\lVert\mathbf{y}-\mathbf{x}\right\rVert^{2}\Bigr{\}},$
and write $\hat{\mathbf{x}}=\operatorname*{arg\,min}_{\mathbf{y}}\{F(\mathbf{y})+\frac{1}{2\lambda}\left\lVert\mathbf{y}-\mathbf{x}\right\rVert^{2}\}$. The convergence criteria is then defined as:
$\nabla F_{\lambda}(\mathbf{x})\coloneqq\lambda^{-1}(\mathbf{x}-\hat{\mathbf{x}}).$
Equivalently, $\nabla F_{\lambda}(\mathbf{x})=\nabla F(\hat{\mathbf{x}})$ in the differentiable case. In other words, the convergence criteria with Moreau envelope uses a surrogate point $\hat{\mathbf{x}}$ instead of the actual iterates of the algorithm. Note that, the convergence criteria using Moreau envelop and proximal gradient mapping is with a constant factor of each other for a $\rho$-weakly convex function [Davis and Drusvyatskiy, 2019]:
$\frac{1}{4}\left\lVert\nabla F_{\nicefrac{{1}}{{2\rho}}}(\mathbf{x})\right\rVert\leq\left\lVert P(\mathbf{x},\nabla f(\mathbf{x}),\rho)\right\rVert\leq\frac{3}{2}\left(1+\frac{1}{\sqrt{2}}\right)\left\lVert\nabla F_{\nicefrac{{1}}{{2\rho}}}(\mathbf{x})\right\rVert.$